From a box with 5 balls numbered from 1 to 5, we draw a ball 500 times with replacement.

Approximate the probability that we will draw the ball with the number “1” less than 80 times.
Answer #1

expected number of times ; ball numbered 1 is drawn =np=500*(1/5)=100

and std devation =sqrt(np(1-p))=8.94

therefore from normal approximation and continuity correction:

probability that we will draw the ball with the number “1” less than 80 times

-P(X<80)=P(Z<(79.5-100)/8.94)=P(Z<-2.29)=0.0110